A 0.050 M concentration of weak base, CH3CH2NH2, was found to ionize at 1.7%. What is the
concentration of CH3CH2NH3⁺ in the solution?
1
Expert's answer
2021-01-04T03:44:11-0500

"CH_3CH_2NH_{2(aq)} + H_2O_{(l)} \u21cc CH_3CH_2NH_{3(aq)}^+ + OH^-_{(aq)}"


"[CH_3CH_2NH_{2}] = 0.050M"


ionization% = "\dfrac{x}{[CH_3CH_2NH_{2}]} × 100\%"


x = [OH-] = "\dfrac{1.7\% × 0.050M }{100\%}" = 8.5 × 10-⁴ M


because the reaction is in equilibrium,

"[CH_3CH_2NH_{3(aq)}^+] = [OH^-]"


"\therefore [CH_3CH_2NH_{3(aq)}^+] = 8.5 × 10^{-4} M"



In conclusion, the concentration of CH3CH2NH3+ in the solution is 8.5 × 10-⁴ M.